OpenStudy (anonymous):

Use the information to solve the triangle. A= 36 degrees, a = 8, and b = 5

hero (hero):

Use law of sines

hero (hero):

\[\frac{\sin A}{a} = \frac{\sin B}{b} = \frac{\sin C}{c}\]

OpenStudy (anonymous):

Do you have any question?